OpenStudy (anonymous):

Evaluate the summation of 2 times negative 2 to the n minus 1 power, from n equals 1 to 7..

OpenStudy (anonymous):

\[\sum_{n=1}^{7} 2(-2)^n\]

OpenStudy (anonymous):

the last n should be ^n-1 not just n

OpenStudy (anonymous):

−256 −84 84 86

OpenStudy (larseighner):

\[\large \sum_{n=1}^{7} 2(-2)^{n-1}\] ?

OpenStudy (anonymous):

yes

OpenStudy (larseighner):

Okay for n=1 the exponent on -2 is 0. So that factor is?

OpenStudy (anonymous):

how do I find the factor

OpenStudy (larseighner):

\[ (-2)^0 = 1\] any number to the zero power is 1 (but not defined for 0 to the 0). So the first term is 2(1).

OpenStudy (larseighner):

So far, the series is 1 + ... Now for the n=2, term.

OpenStudy (anonymous):

ok1 +2 ?

OpenStudy (larseighner):

When n=2, what is: \[\large 2(-2)^{n-1}\]

OpenStudy (anonymous):

=4 -4 ^n-1

OpenStudy (larseighner):

For n=2. \[\large 2(-2)^{2-1} = 2(-2)^1 = 2(-2) = -4\] So the sum so far is 1 + (-4) + ... Now what if n=3?